Betty has 10 more dimes than quarters. If she has $3.45, how many coins does she have?

Answer by stanbon(75887)

Betty has 10 more dimes than quarters. If she has $3.45, how many coins does she have?
---
d = q + 10
10d + 25q = 345 cents
----
Substitute for "d" and solve for "q";
10(q+10) + 25q = 345
35q + 100 = 345
35q = 245
q = 7 (# of quarters)
---
Since d = q+10, d = 17 (# of dimes)
================
Cheers,
Stan H.